# Symbiosis: Multi-Adapter Inference and Fine-Tuning

### v2

@@ Abstract.
Parameter-efficient fine-tuning (PEFT) allows model builders to capture the task-specific parameters into adapters, which are a fraction of the size of the original base model. Popularity of PEFT technique for fine-tuning has led to the creation of a large number of adapters for popular Large Language Models (LLMs). However, existing frameworks fall short in supporting inference or fine-tuning with multiple adapters in the following ways. 1) For fine-tuning, each job needs to deploy its dedicated base model instance, which results in excessive GPU memory consumption and poor GPU utilization. 2) While popular inference platforms can serve multiple PEFT adapters, they do not allow independent resource management or mixing of different PEFT methods. 3) They cannot make effective use of heterogeneous accelerators. 4) They do not provide privacy to users who may not wish to expose their fine-tuned parameters to service providers. In Symbiosis, we address the above problems by enabling the as-a-service deployment of the base model. The base model layers can be shared across multiple inference or fine-tuning processes. Our split-execution technique decouples the execution of client-specific adapters and layers from the frozen base model layers offering them flexibility to manage their resources, to select their fine-tuning method, to achieve their performance goals. Our approach is transparent to models and works out-of-the-box for most models in the transformers library. Our evaluation on Llama2-13B shows that compared to baseline, Symbiosis can fine-tune 4X more adapters on the same set of GPUs in the same amount of time.
conference: ; ;
@@ 1. Introduction
We address the above problems in Symbiosis. Symbiosis shares the common model parameters across the clients for inference and/or fine-tuning jobs. The shared fraction of the model is referred to as base model. We transparently separate the execution of the base model (in base executor) from that of the job specific components, namely, attention and adapter layers, e.g., LoRA, (in client). This split execution offers the following benefits. (1) The separation of execution allows a single base executor to serve multiple inference or fine-tuning clients, where the base executor can be placed in one or more GPUs independently from its clients. This means the base executor can span across multiple GPUs and the clients can be either co-located with the base executor, placed on different GPUs, or even placed on a different node. This is important because the clients’ resource consumption pattern differs from that of the base executor. The client contains request’s runtime state, namely, KV cache or optimizer state, which grows linearly with each new generated token (for inference) or each new pass (for fine-tuning). In contrast, the base executor is stateless. Its resource consumption directly reflects the number of tokens being actively processed at any given moment. Therefore, by separating the base executor from the client, the client can be placed independently and remains insulated from runtime state fluctuations caused by other clients. (2) The client is computationally less intensive, thus they can be hosted on less powerful GPUs or CPUs with little performance degradation. (3) Each user can independently configure its client for an inference or fine-tuning job with a preferred PEFT method while sharing the base model instance. Users opt for different PEFT methods for their different use-case specific strengths, inference overheads, and fine-tuning resource requirements (Lialin et al., 2024; Liu et al., 2022; Nikdan et al., 2024; LeewayHertz, 2025). (4) Different clients using the common base executor need not be in lockstep for the execution of base model layers. Each client can drive its execution at a different rate while opportunistically sharing the execution of the base model with other clients at a layer granularity.
@@
Hardware upgrade with the release of new GPUs is often staggered, which results in nodes in a cluster containing different generations of GPUs. Lack of consideration for their different capabilities degrades the workload performance. Symbiosis proposes a better way to leverage their different capabilities. It decouples the execution of computationally heavy base model from the comparatively lighter clients. This allows the base model to always use the more powerful devices with the clients being potentially placed on less powerful devices, without significant impact on inference or fine-tuning performance. Moreover, Symbiosis also enables execution of clients on CPUs. This is particularly useful for jobs with large KV cache (e.g., with longer sequence lengths) that require 100s of GBs of memory. Longer contexts are common in Retrieval Augmented Generation (RAG) (Lewis et al., 2020), where an additional context is retrieved from a knowledge source and appended to the original prompt. Leveraging the larger CPU memory for such jobs reduces fine-tuning and inference cost. We demonstrate in the evaluation section that for 64K context length, Symbiosis performs CPU-GPU inference and achieves 33% speed up compared to the GPU-only baseline.
@@
With Symbiosis, multiple customers can safely leverage a third-party base model service to fine-tune their models. Even though the popular base model parameters are publicly available, customers may not wish to expose their adapter parameters (which may be trained on customers’ confidential data) and activations to a third-party base model service provider. The challenge of separating the adapter parameters is naturally addressed by Symbiosis’s decoupled execution. This allows a tenant to host the tenant-specific computation in a secure environment. Moreover, Symbiosis also provides a mechanism that avoids exposing the activations that are communicated across client and base executor in order to protect the adapter parameters against model extraction attacks (Liang et al., 2023; Wang et al., 2025; Gong et al., 2020). These attack observe the model activations to infer the model parameters.
@@ 2.1. Model Sharing in Fine Tuning and Inference
Most commonly used PEFT fine-tuning systems (like PyTorch, HF-Trainer (Wolf et al., 2020b)) target single adapter fine-tuning where multi-adapter fine-tuning can only be achieved via multiple isolated tuning jobs. MixLoRA (Li et al., 2024) supports fine-tuning of multiple LoRA adapters for multiple experts in Mixture of Expert model. In contrast, Symbiosis proposes a general purpose system. FlexLLM (Oliaro et al., 2025) enables sharing of the base model across multiple fine-tuning jobs. However, Symbiosis offers better as-a-service model because it isolates the clients’ variable GPU memory demand (from growing KV cache or optimizer state) from each other and from the base executor. We demonstrate that such isolation allows better use of heterogeneous resources and enables privacy of client activations and parameters. mLoRA (Ye et al., 2023; Zhengmao et al., 2023) can simultaneously fine-tune multiple LoRA adapters by sharing the base model across the trainers. However, it only supports variants of LoRA fine-tuning methods. Whereas, Symbiosis allows fine-tuning of wider variety PEFT methods, e.g., IA3, P-tuning, Prefix-tuning etc. Second, unlike mLoRA, Symbiosis does not need to modify the model or user code. Therefore, Symbiosis can work with all the supported models in HuggingFace Transformers library (Wolf et al., 2020b) out-of-the-box. Finally, a key difference between mLoRA and Symbiosis is that in Symbiosis, each client is independent and is a driver of its training or inference. Because of this design principle each client can independently decide its training speed, select different PEFT fine-tuning method, and determine its physical location.
@@
For inference, vLLM (Kwon et al., 2023; Sheng et al., 2024), Punica (Chen et al., 2023), LoRAX (Contributors, 2024), Caraserve (Li et al., [n. d.]) enable sharing of a base model across different adapters and reduce memory consumption. Moreover, their heterogeneous batching allows LoRA adapters with different characteristics, e.g., rank, to be batched together. dLoRA (Wu et al., 2024b) dynamically switches between merged and un-merged adapter to reduce request latency. However, during inference in these systems, different adapters execute as a part the same process as the base model. This means, the client-specific data, i.e., adapter parameters, KV cache and activations, cannot be shielded from being accessed by the base model service provider. Also, clients cannot isolate themselves from the changing resource demands (e.g., increasing KV cache) of other clients. Moreover, the above platforms execute multiple adapters in lock step. Which means the clients with shorter sequences or smaller batches need to wait for the execution of the clients with longer sequences or larger batches at each layer. Symbiosis’s split execution addresses these challenges by isolating the clients from from each other and only opportunistically sharing the execution of base model layers.
@@ 3.1. Design Goals
Work Inference FT Model Sharing Client Independence Flexible Placement Model Transparency Batched Inference-FT Multiple PEFT Adapter Privacy
@@ 3.2. Split Execution in Symbiosis
3.3. Flexible Placement of Clients and Base Executor
@@ 3.5. Client - Base Executor Communication
Figure 7. Per-layer wait time at the base executor for inference clients of Llama2-7B. In local configuration clients and base executor are on the same GPU, whereas for remote configuration the clients are on another GPU.
@@ 3.6. Independent Execution
In Symbiosis, we eliminate the above batching requirement using the following two insights. First, since the layers on the base executor are frozen the parameters are not be updated during a backward pass. Second, for the most popular LLM architectures, (e.g., Llama, GPT, Gemma, Bert), linear and 1D convolution layers constitute the largest fraction of the model parameters. For these layers, the input and output tensors are not involved in gradient calculations. Specifically, the gradient of output w.r.t. input resolves to the parameters themselves. Hence, Symbiosis does not store the input/output tensors for linear and 1D convolution base model layers, even for the fine-tuning request. Instead, it performs a matrix-multiplication between the output gradients and the parameters to generate the required gradients for these layers during the backward pass. This breaks the lockstep, while also providing significant memory savings by not requiring to store input/output tensors for each client at the base executor.
@@ 3.7. Opportunistic Batching
Symbiosis also leverages the insight that for $Conv.1D$ and $nn.Linear$ layer computation, the position of token does not matter in a sequence. Therefore, we are able to flatten all batch-size X sequence-length inputs received from different clients into 1-dimensional sequence of tokens. This allows Symbiosis to avoid padding (which is required to accommodate different size inputs) and save wasted computation.
@@ 3.8. Privacy for Multi-Tenancy
Figure 8. In Symbiosis, the base executor has access to the activations marked in blue colored lines. Therefore, the effect of adapter parameters ($Wa$ $Wb$) can be computed with $(C-B)/A$.
@@
The calculate the effect of noise, we leverage the insight that common LLMs base layers are either $nn.Linear$ or $nn.Conv1D$ layers. For the layers that do not follow linearity, such as $Sigmoid$, $Softmax$, it is not possible to separate the effect of noise from the noisy output and produce the original output. Another challenge with this privacy preserving technique is the presence of bias in $nn.Linear$ or $nn.Conv1D$ layers, which prevents the calculation of equation 2 below. To address this, we also host an alternate execution flow in the base executor which nullifies the effect of bias and returns $n_{effect}$.
@@ 4. Evaluation
For workload, we generate randomly initialized input tensors of desired batch size and sequence lengths. Since, the output with Symbiosis is exactly identical to that of the baseline, the content of the input is not relevant to the performance metrics below. For most fine-tuning experiments, we use a batch size value of 2. This is because the GPU memory consumption is proportional to the product of number of clients, batch size and sequence length. Therefore, smaller batch sizes are best suited to demonstrate the performance with longer sequences and increasing number of clients.
We use LoRA adapters for the following experiments. However, Symbiosis supports other fine-tuning methods such as IA3 and prefix tuning. Table 2 shows the performance of fine-tuning a LoRA adapter with different configurations (Rank, Fine-tuned layers). As compared to the LoRA rank, addition of fine-tuned layers contributes more to increased latency. In the evaluation below, we use the LoRA3 adapter.
@@ 4.1.1. Single Fine-Tuning Job
In Figure 11, we compare the memory consumption of Symbiosis with the baseline for fine-tuning of a single rank-8 LoRA adapter. The Symbiosis without memory optimized backward pass increases the memory requirement compared to baseline. This increase is from having to maintain input and output tensors on client and server side for backward pass. Even though, we don’t use the client side tensors for gradient computation, the client’s computation graph keeps track of the tensors to perform the backward pass. Moreover, each base executor layer also maintains input and output tensors to perform the backward pass. However, with optimized backward pass the memory consumption of the base executor reduces significantly. This is because we do not save the input and output tensors across the layers for gradient computation. This results is almost constant base executor memory footprint.
@@ 4.2.2. Multi GPU Fine Tuning
When measured in terms of memory, for Llama2-13B model, FSDP occupies 17GB of memory on each of the two GPUs. For comparison, this means that 4 FSDP processes can be run in parallel on 2 GPUs (as shown in Figure 17) to fine-tune 4 adapters. In contrast, Symbiosis can fine-tune 8 adapters in half of the time (4X benefit over FSDP).
@@
Figure 17 shows the throughput with increasing clients for Gemma2-27B model. For comparison, the baseline FSDP over 8 GPUs fine-tunes a single adapter at 32 tokens/s with batch size of 2 and sequence length of 64. The primary source of overhead with both baseline and Symbiosis is from parameter fetching. Additionally, the baseline needs to exchange gradients for fine-tuning a common adapter. As a result, Symbiosis outperforms FSDP with 8 adapters by 3X (and even 4). This configuration is equivalent to FSDP, because with FSDP each of 8 shards processes separate tokens.

### v3

@@ Abstract.
Parameter-efficient fine-tuning (PEFT) allows model builders to capture the task-specific parameters into adapters, which are a fraction of the size of the original base model. Popularity of PEFT technique for fine-tuning has led to the creation of a large number of adapters for popular Large Language Models (LLMs). However, existing frameworks fall short in supporting inference or fine-tuning with multiple adapters in the following ways. 1) For fine-tuning, each job needs to deploy its dedicated base model instance, which results in excessive GPU memory consumption and poor GPU utilization. 2) While popular inference platforms can serve multiple PEFT adapters, they do not allow independent resource management or mixing of different PEFT methods. 3) They cannot make effective use of heterogeneous accelerators. 4) They do not provide privacy to users who may not wish to expose their fine-tuned parameters to service providers. In Symbiosis, we address the above problems by enabling the as-a-service deployment of the base model. The base model layers can be shared across multiple inference or fine-tuning processes. Our split-execution technique decouples the execution of client-specific adapters and layers from the frozen base model layers offering them flexibility to manage their resources, to select their fine-tuning method, to achieve their performance goals. Our approach is transparent to models and works out-of-the-box for most models in the transformers library. We demonstrate the use of Symbiosis to simultaneously fine-tune 20 Gemma2-27B LoRA adapters on 8 GPUs.
Model Sharing, PEFT
copyright: acmlicensedcopyright: nonejournalyear: 2025conference: ACM Symposium on Cloud Computing; November 19–21, 2025; Online, booktitle: ACM Symposium on Cloud Computing (SoCC ’25), November 19–21, 2025, Online, doi: 10.1145/3772052.3772230isbn: 979-8-4007-2276-9/25/11
@@ 1. Introduction
We address the above problems in Symbiosis. Symbiosis shares the common model parameters across the clients for inference and/or fine-tuning jobs. The shared fraction of the model is referred to as base model. We transparently separate the execution of the base model (in base executor) from that of the job specific components, namely, attention and adapter layers, e.g., LoRA, (in client). This split execution offers the following benefits. (1) The separation of execution allows a single base executor to serve multiple inference or fine-tuning clients, where the base executor can be placed in one or more GPUs independently from its clients. This means the base executor can span across multiple GPUs and the clients can be either co-located with the base executor, placed on different GPUs, or even placed on a different node. This is important because the clients’ resource consumption pattern differs from that of the base executor. The client contains request’s runtime state, such as, KV cache or optimizer state, which grows linearly with each new generated token (for inference) or each new pass (for fine-tuning). In contrast, the base executor is stateless. Its resource consumption directly reflects the number of tokens being actively processed at any given moment. Therefore, by separating the base executor from the client, the client can be placed independently and remains insulated from runtime state fluctuations caused by other clients. (2) The clients are computationally less intensive, thus they can be hosted on less powerful GPUs or CPUs with little performance degradation. (3) Each user can independently configure its client for an inference or fine-tuning job with a preferred PEFT method while sharing the base model instance. Users opt for different PEFT methods for their different use-case specific strengths, inference overheads, and fine-tuning resource requirements (Lialin et al., 2024; Liu et al., 2022; Nikdan et al., 2024; LeewayHertz, 2025). (4) Different clients using the common base executor need not be in lockstep for the execution of base model layers. Each client can drive its execution at a different rate, while opportunistically sharing the execution of the base model with other clients at a layer granularity.
@@
Hardware upgrade with the release of new GPUs is often staggered, which results in nodes in a cluster containing different generations of GPUs. Lack of consideration for their different capabilities degrades the workload performance. Symbiosis proposes a better way to leverage their different capabilities. It decouples the execution of computationally heavy base model from the comparatively lighter clients. This allows the base model to always use more powerful devices with the clients being potentially placed on less powerful devices, without significant impact on inference or fine-tuning performance. Moreover, Symbiosis also enables execution of clients on CPUs. This is particularly useful for jobs with large KV cache (e.g., with longer sequence lengths) that require 100s of GBs of memory. Longer contexts are common in Retrieval Augmented Generation (RAG) (Lewis et al., 2020), where an additional context is retrieved from a knowledge source and appended to the original prompt. Leveraging the larger CPU memory for such jobs reduces fine-tuning and inference cost. We demonstrate in the evaluation section that for 64K context length, Symbiosis performs CPU-GPU inference and achieves 33% speed up compared to the GPU-only baseline.
@@
With Symbiosis, multiple customers can safely leverage a third-party base model service to fine-tune their models. Even though the popular base model parameters are publicly available, customers may not wish to expose their adapter parameters (which may be trained on customers’ confidential data) and activations to a third-party base model service provider. The challenge of separating the adapter parameters is naturally addressed by Symbiosis’s decoupled execution. This allows a tenant to host the tenant-specific computation in a secure environment. Moreover, Symbiosis also provides a mechanism that avoids exposing the activations that are communicated across client and base executor in order to protect the adapter parameters against model extraction attacks (Liang et al., 2023; Wang et al., 2025; Gong et al., 2020). These attacks observe the model activations to infer the model parameters.
@@ 2.1. Model Sharing in Fine Tuning and Inference
Most commonly used PEFT fine-tuning systems (like PyTorch, HF-Trainer (Wolf et al., 2020b)) target single adapter fine-tuning where multi-adapter fine-tuning can only be achieved via multiple isolated tuning jobs. MixLoRA (Li et al., 2024) supports fine-tuning of multiple LoRA adapters for multiple experts in Mixture of Expert model. In contrast, Symbiosis proposes a general purpose system. FlexLLM (Oliaro et al., 2025) enables sharing of the base model across multiple fine-tuning jobs. However, it relies on static compilation of models (fle, 2025) to generate execution graphs and parallelization strategies based on the model’s architecture. Therefore, any structural changes to model architecture, such as adding new PEFT adapters, cannot occur at runtime. In contrast, Symbiosis is designed for Transformers platform to handle the addition (or removal) of clients (i.e., PEFT adapters) at runtime. This gives tenants runtime control over their clients’ placement, thus enabling better resource isolation and privacy. mLoRA (Ye et al., 2023; Zhengmao et al., 2023) can simultaneously fine-tune multiple LoRA adapters by sharing the base model across the trainers. However, it only supports variants of LoRA fine-tuning methods. Whereas, Symbiosis allows fine-tuning of wider variety PEFT methods, e.g., IA3, P-tuning, Prefix-tuning etc. Second, unlike mLoRA, Symbiosis does not need to modify the model or user code. Therefore, Symbiosis can work with all the supported models in HuggingFace Transformers library (Wolf et al., 2020b) out-of-the-box. Finally, a key difference between mLoRA and Symbiosis is that in Symbiosis, each client is independent and is a driver of its training or inference. Because of this design principle each client can independently decide its training speed or select different PEFT fine-tuning method.
@@
For inference, vLLM (Kwon et al., 2023; Sheng et al., 2024), Punica (Chen et al., 2023), LoRAX (Contributors, 2024), Caraserve (Li et al., [n. d.]) enable sharing of a base model across different adapters and reduce memory consumption. Moreover, their heterogeneous batching allows LoRA adapters with different characteristics, e.g., rank, to be batched together. dLoRA (Wu et al., 2024b) dynamically switches between merged and un-merged adapter to reduce request latency. However, during inference in these systems, different adapters execute as a part the same process as the base model. This means, the client-specific data, i.e., adapter parameters, KV cache and activations, cannot be shielded from being accessed by the base model service provider. Also, clients cannot isolate themselves from the changing resource demands (e.g., increasing KV cache) of other clients. Moreover, the above platforms execute multiple adapters in lock step. Which means the clients with shorter sequences or smaller batches need to wait for the execution of the clients with longer sequences or larger batches at each layer. Symbiosis’s split execution addresses these challenges by isolating the clients from each other and only opportunistically sharing the execution of base model layers.
@@ 3.1. Design Goals
Work Inference FT Model Sharing Client Independence Flexible Placement Model Transparency Batched Inf-FT Multiple PEFT Adapter Privacy
@@ 3.2. Split Execution in Symbiosis
3.3. Flexible Placement
@@ 3.5. Client - Base Executor Communication
Figure 7. Per-layer wait time at the base executor for local and remote inference clients of Llama2-7B.
@@ 3.6. Independent Execution
In Symbiosis, we eliminate the above batching requirement using the following two insights. First, since the layers on the base executor are frozen the parameters are not updated during a backward pass. Second, for the most popular LLM architectures, (e.g., Llama, GPT, Gemma, Bert), linear and 1D convolution layers constitute the largest fraction of the model parameters. For these layers, the input and output tensors are not involved in gradient calculations. Specifically, the gradient of output w.r.t. input resolves to the parameters themselves. Hence, Symbiosis does not store the input/output tensors for linear and 1D convolution base model layers, even for the fine-tuning request. Instead, it performs a matrix-multiplication between the output gradients and the parameters to generate the required gradients for these layers during the backward pass. This breaks the lockstep, while also providing significant memory savings by not requiring to store input/output tensors for each client at the base executor.
@@ 3.7. Opportunistic Batching
Symbiosis also leverages the insight that for $Conv.1D$ and $n$.Linearnn.Linear layer computation, the position of token does not matter in a sequence. Therefore, we are able to flatten all batch-size X sequence-length inputs received from different clients into 1-dimensional sequence of tokens. This allows Symbiosis to avoid padding (which is required to accommodate different size inputs) and save wasted computation.
@@ 3.8. Privacy for Multi-Tenancy
Figure 8. In Symbiosis, the base executor has access to the activations marked in blue colored lines. Therefore, the effect of adapter parameters (WaWa . WbWb) can be computed with (CB)/A(C-B)/A.
@@
The calculate the effect of noise, we leverage the insight that common LLMs base layers are either $n$.Linearnn.Linear or $n$.Conv1Dnn.Conv1D layers. For the layers that do not follow linearity, such as $Sigmoid$, SoftmaxSoftmax, it is not possible to separate the effect of noise from the noisy output and produce the original output. Another challenge with this privacy preserving technique is the presence of bias in $n$.Linearnn.Linear or $n$.Conv1Dnn.Conv1D layers, which prevents the calculation of equation 2 below. To address this, we also host an alternate execution flow in the base executor which nullifies the effect of bias and returns $n_{effect}$.
@@ 4. Evaluation
For workload, we generate randomly initialized input tensors of desired batch size and sequence lengths. Since, the output with Symbiosis is exactly identical to that of the baseline, the content of the input is not relevant to the performance metrics below. For most fine-tuning experiments, we use a batch size value of 2, a common default configuration across popular platforms (aws, 2025; ibm, 2025; uns, 2025). This is because the GPU memory consumption is proportional to the product of number of clients, batch size and sequence length. Therefore, smaller batch sizes are best suited to demonstrate the performance with longer sequences and increasing number of clients.
We use LoRA adapters for the following experiments. However, Symbiosis supports other fine-tuning methods such as IA3 and prefix tuning. Table 2 shows the performance of fine-tuning a LoRA adapter with different configurations (Rank, Fine-tuned layers). As compared to the LoRA rank, addition of fine-tuned layers contributes more to increased latency. Therefore, in the evaluation below, we use the LoRA3 adapter, which uses maximum possible LoRA layers, namely, q, k, v, o.
@@ 4.1.1. Single Fine-Tuning Job
In Figure 11, we compare the memory consumption of Symbiosis with the baseline for fine-tuning of a single rank-8 LoRA adapter. The Symbiosis without memory optimized backward pass increases the memory requirement compared to baseline. This increase is from having to maintain two copies of input and output tensors (on client side and base executor side) for backward pass. Even though, we don’t use the client side tensors for gradient computation, the client’s computation graph keeps track of the tensors to perform the backward pass. Symbiosis’s Memory Optimized (MO) version addresses this problem by eliminating the need of the base executor-side tensor copy, thus making Symbiosis’s memory footprint similar to that of the baseline. Moreover, unlike the baseline, the optimization provides constant base executor-side memory footprint even with increasing sequence length.
@@ 4.2.2. Multi GPU Fine Tuning
When measured in terms of memory, for Llama2-13B model, FSDP occupies 17GB of memory on each of the two GPUs. For comparison, this means that 4 FSDP processes can be run in parallel on 2 GPUs (as shown in Figure 17) to fine-tune 4 adapters. In contrast, Symbiosis can fine-tune 4 adapters in almost half the time.
@@
Figure 17 shows the throughput with increasing clients for Gemma2-27B model. For comparison, the baseline FSDP over 8 GPUs fine-tunes a single adapter at 32 tokens/s with batch size of 2 and sequence length of 64. The primary source of overhead with both baseline and Symbiosis is from parameter fetching. Additionally, the baseline needs to exchange gradients for fine-tuning a common adapter. As a result, Symbiosis with 8 adapters (clients) outperforms a comparable single adapter 8-GPU FSDP baseline, where each shard processes separate tokens.
@@ 4.3.2. Inference with Heterogeneous Client
Configuration Latency
Small & Small 0.30
Small & Large 3.74
Large & Large 6.94
Table 4. vLLM (v0.7.0) prefill response time for batched inference of small and large requests for Llama2-7B model. Small: sequence length=1, Large: sequence length=512.
Prefill Generation
Request Rate Average Throughput Latency Throughput Latency
(reqs/s) Batch Size (tokens/s) (s) (tokens/s) (s)
No Lockstep 0.72 1.1 18912 8.7 13.7 2.3
Opportunistic Batching 0.87 2.14 20392 7.2 17.6 2
Table 5. Average latency and aggregate throughput of 32 inference jobs in 8-GPU sharded local configuration of Llama2-13B. Opportunistic batching achieves 20% higher request rate. Moreover, it improves token throughput while lowering request latency.